Show the method of dimension to verify weather or not the relation s=ut+½at² is correct
1
Expert's answer
2018-12-19T09:40:37-0500

Answer on Question 83857, Physics, Other

Question:

Show the method of dimension to verify whether or not the relation s=ut+½at² is correct.

Answer:

If the relation is dimensionally correct, its left-hand side is equal to the right-hand side. Let's check the dimensions:

"[s] =[displacement] = L,""[u] = [speed] = L\/T,""[t] = [time] = T,""[a] = [acceleration] = L\/T^2,""[t^2] = T^2,""[ut + 1\/2at^2] = (L\/T)*L + (L\/T^2)*T^2 = L + L = L,""L = L."

As we can see, the relation s=ut+½at² is dimensionally correct.